OpenStudy (anonymous):

Which shows the fractions in increasing order? four-ninths, five-ninths, five-twelfths, seven-twelfths four-ninths, five-twelfths, seven-twelfths, five-ninths five-twelfths, four-ninths, five-ninths, seven-twelfths seven-twelfths, five-ninths, five-twelfths, four-ninths

OpenStudy (anonymous):

First it'll be easier to see if you write them all out numerically, Your four fractions are \[ \frac{4}{9}, \frac{5}{12},\frac{7}{12}, \frac{5}{9} \] to make it even easier, you probably want to make sure all of the denometers are the same. What's the smallest number to 9 and 12 both go into?

OpenStudy (anonymous):

4*9 = 3*12 = 36 \[ \frac{4}{4}\cdot\frac{4}{9} \ \ , \ \ \frac{3}{3}\cdot\frac{5}{12} \ \ \ , \ \ \ \frac{3}{3}\cdot\frac{7}{12}, \ \ \ \frac{4}{4} \cdot \frac{5}{9} \] = \[ \frac{16}{36}, \ \ \frac{15}{36}, \ \ \frac{21}{36}, \ \ \frac{20}{36}\]

OpenStudy (anonymous):

ty